OpenStudy (anonymous):

please help! factor the trinomial. v^2-17v+70

OpenStudy (akshay_budhkar):

v^2-7v-10v+70 =v(v-7)-10(v-7) =(v-7)(v-10)

OpenStudy (anonymous):

You're looking for numbers a and b with a+b=-17 and a*b=70 You could see that a=-10 and b=-7 work. So we get (v+a)(v+b)=(v-10)(v-7)

OpenStudy (anonymous):

ok thanks